OpenStudy (anonymous):

Please help -show work.7 7/9 *1 11/45 Im really tired i want to go to bed can any one help :(

OpenStudy (anonymous):

7 7/9 = 70/9 1 11/45 = 56/45 So the product is : (70x56)/(9x45) = 3920 / 405 = 784 / 81

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

can you help me with this problem.... 8 4/15 *(-2/5)

OpenStudy (anonymous):

8 4/15 = 124 / 15 So the product is = (124x-2) / (15x5) = -248 / 75

OpenStudy (anonymous):

# 24. 34 *(-4/5)

OpenStudy (anonymous):

i don't get this can you do this one?